Shift report 20000901 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Sep 01 2000 - 07:56:06 EDT

  • Next message: Yury Blyakhman: "Re: Thanks,. yury!"

    ------------------------------------------------------------
    Date: Fri Sep  1 07:56:05 2000, Shift: 00:00-08:00
    Supervisor: HI, with: AJ
    Log book#5 pages 70 - 71
    
    Shift Summary:
    ------------------------------------------------------------
    1.  No beam from 0:20-3:51 due to quench link interlock.
    2.  Run 2492 with FF at 12 deg, BF 4 deg and MIDS at 40 deg
    Run 2493 with FF at 4 deg, BF 4 deg and MIDS at 40 deg
    3.  BF platform were not able to move yet.
    One Tile (Tile5_3) did not have a signal initially but fixed.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Sep 01 2000 - 07:56:47 EDT